I am a 2-digit number. The sum of my digits is 10. The tens and ones digits are odd. The tens digit is greater than the ones digit. What numbers can I be?
Answer by jim_thompson5910(35256)

Odd numbers that add to 10:
1+9=10, 3+7=10, 5+5=10, 7+3=10, 9+1=10

Since the tens digit is greater than the ones digit, we only have the two numbers: 73 and 91